OpenStudy (anonymous):

Simplify the expression 18x2/3x6 .

OpenStudy (anonymous):

72

OpenStudy (anonymous):

does "simplify" mean compute the number?

OpenStudy (anonymous):

yes simplify means solve or put in simplest form if variables are there

OpenStudy (anonymous):

\[\frac{18\times 2}{3\times 6}=\frac{18\times 2}{18}=2\]

OpenStudy (anonymous):

thanks(: